The number of accidents in a production facility has a Poisson distribution with a mean of 2.7 per month.For a given month what is the probability that there will be more than three accidents?


1
Expert's answer
2021-05-07T12:15:37-0400

Given mean "\lambda=2.7,"


Let "X" denote the number of accidents-


 Probability that there will be more than three accidents-

"P(X>3)=1-[P(X=0)+P(X=1)+P(X=2)+P(X=3)]"

"=1-[\dfrac{e^{-\lambda}.\lambda^0}{0!}+\dfrac{e^{-\lambda}.\lambda^1}{1!}+\dfrac{e^{-\lambda}.\lambda^2}{2!}+\dfrac{e^{-\lambda}.\lambda^3}{3!}]"


"=1-[\dfrac{e^{-2.7}.(2.7)^0}{1}+\dfrac{e^{-2.7}.(2.7)^1}{1}+\dfrac{e^{-2.7}.(2.7)^2}{2}+\dfrac{e^{-2.7}.(2.7)^2}{6}]\\[9pt]=1-(0.0672+0.08145+0.24496+0.22046)\\[9pt]=1-0.61767\\[9pt]=0.38233"